### v1

# NeUQI: Near-Optimal Uniform Quantization Parameter Initialization

Abstract
Large language models (LLMs) achieve impressive performance across domains but face significant challenges when deployed on consumer-grade GPUs or personal devices such as laptops, due to high memory consumption and inference costs. Post-training quantization (PTQ) of LLMs offers a promising solution that reduces their memory footprint and decoding latency. In practice, PTQ with uniform quantization representation is favored for its efficiency and ease of deployment since uniform quantization is widely supported by mainstream hardware and software libraries. Recent studies on $\geq 2$-bit uniform quantization have led to noticeable improvements in post-quantization model performance; however, they primarily focus on quantization methodologies, while the initialization of quantization parameters is underexplored and still relies on the suboptimal Min-Max strategies. In this work, we propose NeUQI, a method devoted to efficiently determining near-optimal initial parameters for uniform quantization. NeUQI is orthogonal to prior quantization methodologies and can seamlessly integrate with them. The experiments with the LLaMA and Qwen families on various tasks demonstrate that our NeUQI consistently outperforms existing methods. Furthermore, when combined with a lightweight distillation strategy, NeUQI can achieve superior performance to PV-tuning, a much more resource-intensive approach.
1 Introduction
In recent years, large language models (LLMs) like ChatGPT [1] have rapidly emerged, demonstrating strong capabilities across various tasks, including open-ended writing, knowledge-based question answering, and code generation. Given the high API costs of proprietary LLMs and the growing performance of open-source alternatives like LLaMA [28, 10] and Qwen [32] families, there is a rising preference for deploying open-source LLMs locally. However, the deployment of large-scale models (e.g., LLaMA 3 70B) is often limited by compute resources and inference efficiency, especially on personal devices or consumer-grade GPUs like a single RTX 4090 with 24GB of memory. In this context, post-training quantization (PTQ) [15], particularly using a uniform quantization scheme, offers a practical solution by converting model weights from bfloat16 to low-bit-width integer formats such as int4/3/2, significantly reducing memory usage and inference latency.
Prior work on uniform PTQ mainly focuses on two directions. The first involves weight and activation processing to reduce quantization difficulty, such as smoothing via rotation transformations [5, 29, 3], integrating rotations into model architectures [21, 13], and reducing weight magnitude [34]. The second emphasizes block-wise and global fine-tuning, including multi-block training [7], rotation-based fine-tuning [21, 13], and direct optimization over quantized weights [22].
Although these studies have made notable progress, they still adopt the conventional Min-Max initialization, which merely determines quantization parameters solely based on the extreme values in the weights while ignoring weight distribution and importance. As revealed in LeanQuant [35], such initialization leads to a mismatch with the quantization objective in the commonly adopted GPTQ framework [8]. Even under ideal conditions, specifically a uniform distribution, Min-Max remains suboptimal, as analyzed in Appendix A. However, little research has focused on improving quantization initialization. Only recently has LeanQuant made a preliminary attempt by introducing a loss-aware strategy. But it still inherits a key limitation of Min-Max: restricting the zero-point to integers, which reduces the flexibility of uniform quantization and results in suboptimal initialization.
To address this gap, we propose NeUQI, a novel initialization method for uniform quantization, which removes the integer constraint on the zero-point, thus enabling a more flexible parameter space. We first apply a diagonal approximation to the quantization reconstruction objective and then decompose it into two tractable subproblems: (1) determining the optimal zero-point under a fixed scale and (2) finding the optimal scale given its corresponding zero-point. Both are efficiently solvable by exploiting the objective’s structure. More importantly, NeUQI is plug-and-play and orthogonal to quantization methods used in the calibration and fine-tuning stages. When combined with our NeUQI, the performance of quantized models significantly improves and even surpasses non-quantized models of comparable memory usage, as illustrated in Figure 1. Furthermore, only applying a lightweight knowledge distillation, NeUQI can outperform PV-tuning [22]—a more sophisticated and resource-intensive fine-tuning method. It underscores the critical role of proper initialization: poor initial conditions are difficult to fix even with heavy fine-tuning, while a well-initialized model can achieve better performance with significantly less effort.
Figure 1: Perplexity on C4 (left) and average accuracy across five common benchmarks (right) with Qwen 2.5 family, plotted against the base-2 logarithm ($\log_{2}$) of model memory usage on the x-axis. The results include the non-quantized models (BF16) and the quantized models using commonly-used GPTQ, or our proposed NeUQI (both 3-bit). More details and results are shown in Section 5.
The main contributions of our work are summarized as follows:
  1. 1.
    We revisit the conventionally adopted Min-Max quantization initialization and identify its fundamental limitation: the restriction of the zero-point to integers, which leads to reduced flexibility and suboptimal initialization.
  2. 2.
    We propose NeUQI, a novel initialization method for uniform quantization that relaxes the zero-point to a continuous variable and exploits the structure of the optimization landscape. NeUQI can be directly combined with existing quantization methods to enhance their performance.
  3. 3.
    Experiments with LLaMA and Qwen families of different sizes on various tasks demonstrate that NeUQI consistently outperforms existing methods. When further combined with a lightweight distillation strategy, NeUQI even surpasses relatively more resource-intensive PV-tuning.
2 Related Work
Uniform Quantization
In the 2-bit to 4-bit quantization regime, prior studies have explored many post-training quantization techniques to mitigate accuracy degradation. Early methods like AWQ [19] and SmoothQuant [31] applied mathematically equivalent transformations using diagonal matrices on activations and their inverses on weights, balancing quantization difficulty and laying the groundwork for later approaches. Subsequent methods generalized the diagonal matrix to invertible forms for better smoothing, though activation-side transformations require online computation, leading to efficiency-focused designs. For example, QuIP [5] uses Kronecker products of two small rotations; DuQuant [18] adopts block-diagonal matrices with repeated rotation; QuIP# [29] and QuaRot [3] use Hadamard transforms; FrameQuant [2] applies a Fourier-based transformation variant. SpinQuant [21] and OSTQuant [13] introduce learnable rotations integrated into the model to avoid runtime overhead. Alternatively, MagR [34] optimizes weights while preserving model functionality to reduce quantization difficulty. In addition, there are fine-tuning-based approaches such as CBQ [7], which performs block-wise tuning, and PV-tuning [22], which jointly optimizes continuous and discrete parameters.
Quantization Parameter Initialization
Most uniform quantization methods rely on the conventional Min-Max initialization, which determines quantization parameters by the extreme weight values. Although Min-Max is straightforward and efficient, this heuristic neglects the impact of initialization on the final model performance. Some methods, such as MagR and OmniQuant [27], introduce shrink-scale variants of Min-Max to suppress outliers and improve precision, but they remain simple extensions of the original method. LeanQuant [35] makes a significant step forward by proposing an optimization-based initialization with grid search. However, its optimization relies on enumerating different Min-Max value pairs, from which the scale and zero-point are computed based on the standard Min-Max formula. As a result, it inherits the core limitations of Min-Max, particularly the constraint of zero-points to integers. In contrast, our NeUQI is proposed to overcome the limitations of the Min-Max paradigm by removing the zero-point constraint and directly optimizing a well-founded initialization objective.
3 Background
Uniform Quantization
Model quantization is, in essence, the process of mapping floating-point parameters to a finite set of quantized values, allowing each parameter to be represented by a low-bit-width index, thereby reducing memory footprint and memory access latency. In the case of uniform quantization, the finite set $\mathcal{Q}$ is defined by three parameters: the scale $s$, the zero-point $z$, and the bit-width $k$. Formally, this set is given by:
$\mathcal{Q}=\left\{q_{i}=s\cdot(z+i)\,\middle|\,i\in\{0,1,\ldots,2^{k}-1\}\right\},$ (1)
where $i$ is the index of quantized values.
To reduce quantization error, each floating-point value is typically mapped to the nearest value in $\mathcal{Q}$. Given a floating-point input $x$, and assuming the parameters $s$, $z$, and $k$ of $\mathcal{Q}$ are known, the uniform quantization function $Q$ is defined as:
$Q_{s,z}(x)=s\cdot\left(\text{clamp}(\lfloor x/s-z\rceil,\,0,\,2^{k}-1)+z\right),$ (2)
[…]
Min-Max Initialization
Min-Max is a conventional initialization method that remains widely used in uniform quantization. Min-Max determines the quantization parameters by mapping the minimum and maximum of the parameter to $s\cdot z$ and $s\cdot(z+2^{k}-1)$, which correspond to the quantized values at indices 00 and $12^{k}-$ 1 in the set $\mathcal{Q}$. Based on the above extremal-value mapping, the scale factor $s$ and the zero point $z$ are directly given by:
$s=\frac{\text{max}(x)-\text{min}(x)}{2^{k}-1},\quad z=\left\lfloor\frac{\text{min}(x)}{s}\right\rceil.$ (3)
Here, $z$ is rounded to the nearest integer. This is because, under the common assumptions that $\text{max}(x)\geq 0$, $\text{min}(x)\leq 0$, and $s>0$, we have $s\geq-\text{min}(x)/(2^{k}-1)$ and $\text{min}(x)/s\leq 0$, which implies $\text{min}(x)/s\in[-(2^{k}-1),0]$. Therefore, after rounding, it follows that $z\in[-(2^{k}-1),0]$, and thus $-z$ lies within the range $[0,2^{k}-1]$, making it representable as a $k$-bit unsigned integer.
4 Method
In this section, we present NeUQI (Near-Optimal Uniform Quantization Initialization), a novel method designed to improve the parameter initialization of uniform quantization. Unlike the conventional Min-Max initialization, which directly determines the scale and zero-point from the observed minimum and maximum values of the weights, NeUQI formulates the initialization as an optimization problem to find near-optimal quantization parameters. We begin by revisiting the quantization loss formulation in Section 4.1, which lays the foundation for our initialization strategy. Section 4.2 then introduces the NeUQI algorithm, detailing its formulation and practical implementation. Finally, in Section 4.3, we introduce a lightweight distillation technique to demonstrate the efficiency of our method, which can further improve the performance of the quantized model with minimal overhead.
4.1 Quantization Loss
Following GPTQ [8], the quantization loss function for a single linear layer is given by
$\mathcal{L}(s,z)$ $=\|XW^{\top}-XQ_{s,z}(W)^{\top}\|_{F}^{2}$ (4)
$=\operatorname{tr}((W-Q_{s,z}(W))H(W-Q_{s,z}(W))^{\top})$ (5)
$=\sum_{i}(W_{i}-Q_{s,z}(W_{i}))H(W_{i}-Q_{s,z}(W_{i}))^{\top},$ (6)
where $X\in\mathbb{R}^{T\times\mathrm{In}}$ denotes the input activations, $W\in\mathbb{R}^{\mathrm{Out}\times\mathrm{In}}$ is the weight matrix of the linear layer, with $\mathrm{In}$ and $\mathrm{Out}$ representing the input and output dimensions respectively. $W_{i}\in\mathbb{R}^{1\times\mathrm{In}}$ denotes the $i$-th row of $W$, and $H=X^{\top}X\in\mathbb{R}^{\mathrm{In}\times\mathrm{In}}$ is commonly referred to as the Hessian matrix. This formulation naturally suggests adopting the row-wise quantization strategy, which aligns with common practices that treat one row as one or multiple groups for quantization.
4.2 NeUQI
Although the full Hessian is efficiently computable for a single linear layer, incorporating it into the optimization is nontrivial. To address this, we adopt a diagonal approximation, which is a standard technique in the machine learning literature [17, 14]. This approximation removes cross-weight interactions and results in a structurally simpler loss function that, despite being globally coupled via shared quantization parameters, is significantly easier to optimize. Under this approximation, the loss function reduces to:
$\mathcal{L}(s,z)=\sum_{i}H_{ii}(Q_{s,z}(w_{i})-w_{i})^{2}$ (7)
To solve this problem efficiently, we decompose the optimization into two sequential steps. First, we fix the scale $s$ and optimize the zero-point $z$, as discussed in Section 4.2.1. Then, based on the solution of the first step, we further optimize the scale $s$, as described in Section 4.2.2.
4.2.1 Optimization of Zero-Point
Following this decomposition, the loss function is further transformed into:
$\mathcal{L}(z)=s^{2}\sum_{i=1}^{n}h_{i}\left(\text{clamp}\left(\lfloor x_{i}-z\rceil,0,2^{k}-1\right)-(x_{i}-z)\right)^{2},$ (8)
where $x_{i}=w_{i}/s$, and $h_{i}=H_{ii}$.
Algorithm 1 Optimal Zero-Point Search over Piecewise Quadratic Function
1:Samples $\{x_{i}\}_{i=1}^{n}$, bit-width $k$
2:Optimal zero-point $z^{*}$ minimizing simplified quantization loss
3:Initialize list of transition points: $\mathcal{T}\leftarrow[]$
4:for each sample $x_{i}$ do
5:     for $j=0$ to $22^{k}-$ 2 do
6:         Compute transition point: $t\leftarrow x_{i}-(j+0.5)$
7:         Compute loss increment: $\Delta\mathcal{L}\leftarrow h_{i}((j-(x_{i}-z))^{2}-(j+1-(x_{i}-z))^{2})$
8:         Add $(t,\Delta\mathcal{L})$ to $\mathcal{T}$
9:     end for
10:end for
11:Sort $\mathcal{T}$ by transition point $t$
12:Initialize coefficients $Az^{2}+Bz+C\leftarrow\sum_{i}h_{i}(2^{k}-1-(x_{i}-z))^{2}$
13:Let $(t_{\text{first}},\delta_{\text{first}})\leftarrow$ the first element of $\mathcal{T}$
14:Compute minimum: $z^{\prime}\leftarrow-B/(2A)$
15:Clamp: $z^{\prime}\leftarrow\text{min}(z^{\prime},t_{\text{first}})$
16:Evaluate loss at $z^{\prime}$, record as current best: $z^{*}\leftarrow z^{\prime}$
17:for each $(t,\delta)$ in $\mathcal{T}$ do
18:     Update $Az^{2}+Bz+C$ using $\delta$
19:     Let next transition point be $t^{\prime}$ (or +$\infty$+ ∞ if none)
20:     Compute minimum: $z^{\prime}\leftarrow-B/(2A)$
21:     Clamp: $z^{\prime}\leftarrow\text{min}(\text{max}(z^{\prime},t),t^{\prime})$
22:     Evaluate loss at $z^{\prime}$, update $z^{*}$ if smaller
23:end for
24:return $z^{*}$ with minimal loss
We first can observe that the single term $\text{clamp}(\lfloor x_{i}-z\rceil,0,2^{k}-1)$ is a piecewise constant function of $z$. The transition points of this function occur at $z=x_{i}-(i+0.5)$, $i\in\{0,1,\cdots,2^{k}-2\}$, which divides the real line into $2^{k}$ intervals. Now consider the single-term quantization error: $(\mathrm{clamp}(\lfloor x_{i}-z\rceil,0,2^{k}-1)-(x_{i}-z))^{2}$, which is a piecewise quadratic function of $z$, also divided into $2^{k}$ intervals. Consequently, the total loss function (aggregated over $n$ samples) consists of $O(n\cdot 2^{k})$ intervals. To find the minimum of the piecewise quadratic function, we first compute the local minimum in each interval and then select the smallest among them. This idea reduces the problem to computing the quadratic, linear, and constant coefficients of the piecewise quadratic loss function in each interval. These coefficients can be efficiently calculated by sorting all transition points and accumulating the corresponding increments, a strategy we refer to as the increment-prefix-sum technique. Specifically, when $\text{clamp}(\lfloor x_{i}-z\rceil,0,2^{k}-1)$ changes from $a$ to $b$, the resulting increment in the loss function is $(b-(x_{i}-z))^{2}-(a-(x_{i}-z))^{2}$. The detailed procedure is described in Algorithm 1.
As previously discussed, the total number of transition points is $O(n\cdot 2^{k})$. Since the coefficient computation involves sorting these points, the overall time complexity is $O(n\cdot 2^{k}\cdot\log(n\cdot 2^{k}))$, which increases overhead substantially as the bit-width $k$ grows. To alleviate this computational burden, we propose a two-stage efficient approximation strategy, which reduces the overall time complexity to $O(n\log n)$:
  • Rough Estimation: At this stage, we approximate $x_{i}-z$ within the low-loss interval $[-\frac{1}{2},2^{k}-\frac{1}{2}]$ by its maximum value $(\frac{1}{2})^{2}$, while leaving the quantization loss unchanged outside this interval, resulting in a smoothed objective:
    $\text{max}\left(0.25,\left(\text{clamp}\left(\lfloor x_{i}-z\rceil,0,2^{k}-1\right)-(x_{i}-z)\right)^{2}\right),$
    Under this approximation, each sample reduces to exactly two transition points, located at $x_{i}-(2^{k}-0.5)$ and $x_{i}+0.5$. Leveraging the same increment-prefix-sum technique, we can determine an approximate zero-point $z_{\mathrm{rough}}$ in $O(n\log n)$ time. The detailed procedure is described in Algorithm 2.
  • Local Search: The original, non-approximated loss function is then evaluated within a narrow interval $[z_{\mathrm{rough}}-1,z_{\mathrm{rough}}+1]$. As the number of transition points in this interval remains limited to at most two per sample, this local search step also maintains a time complexity of $O(n\log n)$. The detailed procedure is described in Algorithm 3.
Figure 2: Line plots of $\mathcal{L}(s,z^{*}(s))$ versus $s$ for the query projection matrix in the 15th transformer block. The horizontal axis shows the ratio of $s$ to the baseline value $(\text{max}(x)-\text{min}(x))/(2^{k}-1)$. The left, center, and right plots correspond to the 1st, 1203rd, and 3709th rows of the query projection matrix. Local minima are indicated with red star ().
4.2.2 Optimization of Scale
After determining the optimal zero-point $z^{*}(s)$ for a given scale $s$ using the above algorithm, the loss function can be simplified into a univariate function that depends only on $s$. Although the loss for a portion of the rows in the weight matrix appears unimodal, many other rows exhibit non-unimodal behavior, as illustrated in Figure 2. Specifically, as shown in the figure, the loss landscape for the 1st row has a single minimum, indicating unimodal behavior, whereas the 1203rd and 3709th rows each exhibit two local minima, clearly showing non-unimodal patterns. As a result, it is challenging to apply continuous optimization techniques like gradient descent or even one-dimensional search methods like ternary search [^1] https://en.wikipedia.org/wiki/Ternary_search for selecting the scale $s$.
To address this issue, we perform a uniformly spaced grid search in the parameter space to ensure sufficient scale exploration. Specifically, the grid points for candidate scales are defined as follows:
$\mathcal{G}_{T}=\left\{\frac{\text{max}(x)-\text{min}(x)}{2^{k}-1}\cdot\frac{i}{T}\,\middle|\,i\in\{1,2,\cdots,T-1,T\}\right\},$ (9)
where $T$ denotes the total number of grid points. We evaluate the quantization error at each grid point and select the scale with the minimum error as the optimal solution.
To further enhance the search efficiency, we propose a coarse-to-fine grid search strategy, which consists of the following two phases:
  • Coarse Grid Search: A preliminary search is conducted using a coarse grid $\mathcal{G}_{T_{c}}$, allowing rapid localization of a promising local region where the optimal solution may reside.
  • Fine Grid Search: After obtaining the coarse search optimal solution $s_{c}$, we construct a local candidate set by selecting the $\lfloor T/(2T_{c})\rfloor$ nearest neighbors on each side of $s_{c}$ from a finer grid $\mathcal{G}_{T}$, along with $s_{c}$ itself. This results in a total of $12\lfloor T/(2T_{c})\rfloor+$ 1 scales, enabling more accurate estimation with relatively low overhead.
This strategy effectively reduces the original $T$ expensive error evaluations—each involving the search for the minimum loss and its corresponding zero-point—to approximately $T_{c}+T/T_{c}$. In particular, when $T_{c}$ is set approximately to $\sqrt{T}$, the total number of evaluations further decreases to around $2\sqrt{T}$. As a result, this approach significantly reduces computational overhead while still providing a good approximation to the optimal solution, achieving a favorable trade-off between quantization accuracy and search efficiency.
4.3 Lightweight Distillation
To better demonstrate NeUQI’s effectiveness, we incorporate additional lightweight distillation experiments for comparison. Specifically, to enhance the alignment between the behaviors of the quantized model and the original non-quantized model, we perform knowledge distillation on a small set of samples. During distillation, only the continuous parameters of the quantized model—specifically, the scales, zero-points, and the RMSNorm layers—are updated. The language model head and embedding layer, which remain in their original precision, are excluded from optimization to avoid distillation-induced bias and reduce memory overhead. The distillation loss is the mean squared error between the final-layer hidden states of the original and quantized models.
5 Experiment
5.1 Baselines and Evaluation
We experiment with our proposed NeUQI on three commonly-used LLM families, covering different sizes: LLaMA 2 (7B, 13B, 70B) [28], LLaMA 3 (8B, 70B) [10] and Qwen 2.5 (7B, 14B, 32B, 72B) [32]. Under uniform quantization, we compare against several representative weight-only post-training quantization methods, including GPTQ [8], MagR [34], PV-tuning [22], LeanQuant [35], as well as two initialization strategy baselines, for different analytical purposes. Following the previous work, all the quantized models are evaluated by measuring perplexity on the WikiText2 (Wiki2[23] and C4 [25] validation sets, and zero-shot accuracy on five benchmarks: ARC-easy (ArcE), ARC-challenge (ArcC[6], PiQA [4], HellaSwag (HellaS[33], and WinoGrande (WinoG[26].
Table 1: Perplexity on Wiki2 and C4, along with zero-shot accuracy on five benchmarks (averaged as Acc) of different quantized models with channel-wise quantization. Arrows / indicate whether the higher or lower result is better, and Size denotes the original model size before quantization.
Size Bits Method Wiki2↓ C4↓ ArcC↑ ArcE↑ HellaS↑ PiQA↑ WinoG↑ Acc↑
LLaMA 2 Family
7B 2 GPTQ 6953 2592 21.93 25.63 25.89 52.23 49.72 35.08
MagR 129 47.55 21.42 34.97 32.29 58.27 50.75 39.54
NeUQI 17.14 17.50 23.98 51.73 36.04 65.89 58.56 47.24
13B 2 GPTQ 1735 433 21.76 26.52 25.85 52.56 51.30 35.60
MagR 47.98 27.94 23.21 26.68 36.70 52.99 54.14 38.74
NeUQI 13.72 14.39 26.19 55.18 37.37 65.94 59.12 48.76
70B 2 GPTQ 60.29 46.11 19.97 28.07 27.05 53.75 50.04 35.78
MagR 68.62 13.95 39.59 71.21 52.81 76.33 69.06 61.80
NeUQI 7.03 8.88 44.71 75.97 53.62 77.64 73.72 65.13
LLaMA 3 Family
8B 2 GPTQ >1e4 >1e4 20.90 23.91 26.04 53.48 48.86 34.64
MagR 387 140 18.43 30.09 27.60 55.98 50.12 36.45
NeUQI 64.47 39.41 24.83 52.10 37.50 63.00 59.04 47.30
70B 2 GPTQ >1e4 >1e4 20.73 25.76 25.76 52.45 48.46 34.63
MagR >1e4 >1e4 21.59 25.59 25.51 52.99 48.46 34.83
NeUQI 56.21 42.06 25.68 53.24 36.98 61.04 55.49 46.49
Qwen 2.5 Family
7B 2 GPTQ 2332 719 21.42 26.52 25.70 50.92 48.93 34.70
MagR 25.38 24.83 23.12 42.09 37.76 60.99 51.30 43.05
NeUQI 37.46 28.77 22.10 47.85 39.56 60.99 58.48 45.80
14B 2 GPTQ 3852 1056 23.12 25.42 25.90 50.92 51.93 35.46
MagR 18.36 19.36 23.21 44.74 38.06 63.44 52.88 44.47
NeUQI 23.58 18.38 37.88 69.32 46.86 72.31 67.40 58.76
32B 2 GPTQ 308 141 21.42 26.09 25.11 51.20 49.41 34.65
MagR 13.16 14.72 25.94 48.91 45.25 68.06 56.51 48.93
NeUQI 18.17 15.77 41.55 71.63 53.38 76.01 72.77 63.07
72B 2 GPTQ 1127 281 22.95 25.00 25.46 51.31 48.70 34.68
MagR 19.25 13.83 31.91 59.64 51.13 72.74 59.59 55.00
NeUQI 10.79 11.36 48.04 78.11 56.18 78.56 75.14 67.21
5.2 Implementation Settings
We evaluate both channel-wise and group-wise quantization at the 2-bit, 3-bit, and 4-bit quantization, where the group size is set to a common configuration of 128. Across all experiments, GPTQ is employed as the unified weight transformation method, ensuring consistent quantization quality and fair comparisons among different approaches. During calibration, 128 samples are drawn from the C4 dataset. Each sample contains 2048 tokens for LLaMA 2, and 4096 tokens for LLaMA 3 and Qwen 2.5, following the previous work. For distillation, 256 samples from the same dataset are used for a single epoch. Moreover, the grid search over the scale hyperparameter is conducted with $T=2048$ and $T_{c}=64$ in our NeUQI method. More implementation details are included in Appendix D.
5.3 Main Results
Under the most challenging 2-bit channel-wise quantization setting, our NeUQI achieves significant improvements over GPTQ and MagR, as shown in Table 1. Although the recent MagR method delivers acceptable performance on the Qwen 2.5 family, especially on perplexity, it performs poorly on the LLaMA family. In contrast, NeUQI consistently demonstrates strong performance across different model architectures and sizes. Similar conclusions hold under the 2-bit group-wise quantization setting, with detailed results shown in Table 6.
Moreover, as the bit-width increases, the performance gap between different methods gradually narrows, as shown in Table 7, Table 8, and Table 9. Our NeUQI continues to show stable advantages under the 3-bit setting, maintaining robustness across architectures. As for 4-bit, where all methods closely match the original non-quantized model, further improvements are limited. Nevertheless, NeUQI remains competitive and exhibits slight advantages in certain cases.
5.4 Discussion
Table 2: Results of different initialization methods based on LLaMA 2 family and LLaMA 3 8B with channel-wise quantization. $\dagger$ denotes the results directly extracted from the original paper.
Size Bits Method Wiki2↓ C4↓ ArcC↑ ArcE↑ HellaS↑ PiQA↑ WinoG↑ Acc↑
LLaMA 2 Family
7B 2 Min-Max+ 498 136 21.67 27.36 26.92 54.24 50.83 36.20
LeanQuant 25.69 27.11 20.99 41.08 31.94 61.64 56.51 42.43
[…]
NeUQI 17.14 17.50 23.98 51.73 36.04 65.89 58.56 47.24
13B 2 Min-Max+ 32.19 25.96 22.61 32.03 32.97 57.40 50.83 39.17
LeanQuant 24.43 20.92 24.32 50.88 38.01 67.19 56.91 47.46
[…]
NeUQI 13.72 14.39 26.19 55.18 37.37 65.94 59.12 48.76
70B 2 Min-Max+ 13.33 14.08 25.94 45.62 38.30 65.56 59.59 47.00
LeanQuant 7.92 10.84 - - - - - -
[…]
NeUQI 7.03 8.88 44.71 75.97 53.62 77.64 73.72 65.13
LLaMA 3 8B
8B 2 Min-Max+ 3016 716 20.99 25.04 26.28 53.16 50.43 35.18
LeanQuant - - 18.26 35.06 31.43 59.30 51.85 39.18
[…]
NeUQI 64.47 39.41 24.83 52.10 37.50 63.00 59.04 47.30
Integer Constraint
[…]
Distillation
As shown in Table 3, augmented with the lightweight distillation training that only uses about 0.5M tokens, NeUQI achieves significantly better results than original PV-tuning. This advantage is more notable considering that PV-tuning requires more memory and incurs longer per-step runtime than standard fine-tuning of the original model, while also demanding substantially more data. When constrained to the same data volume, its performance degrades markedly, falling even below that of the standard GPTQ method. This reveals the importance of proper initialization, since poor initialization seems difficult to fix even with heavy fine-tuning.
Table 3: Results of distillation based on LLaMA 2 7B with 2-bit quantization and group size of 128. $\dagger$ denotes the results directly extracted from the original paper, and Tokens denotes the total number of training tokens.
Method Tokens Wiki2↓ C4↓ ArcC↑ ArcE↑ HellaS↑ PiQA↑ WinoG↑ Acc↑
GPTQ ~0.5M 14.95 13.31 25.34 49.20 39.13 67.63 58.96 48.05
PV-tuning ~0.5M 14.09 13.95 23.21 43.35 37.38 64.80 58.17 45.38
PV-tuning ~1B 8.49 10.78 - - - - - 52.17
NeUQI ~0.5M 8.38 9.81 32.00 66.79 47.73 73.34 64.01 56.77
Comparison with Non-Quantized Models
With comparable memory usages, the quantized models using our NeUQI consistently outperform the non-quantized ones, as shown in Table 4. Although limited by computational resources and unable to experiment on larger models, these results demonstrate the potential of NeUQI-based low-bit-width quantization when applied to ultra-large models (e.g., DeepsSeek V3 671B [20]). They often lack smaller versions and remain difficult to deploy, particularly in resource-limited environments, even with 8-bit or 4-bit quantization.
Table 4: Performance comparison between quantized and non-quantized models based on the Qwen 2.5 family with comparable main body memory usages.
Method Memory Size Bits Wiki2↓ C4↓ ArcC↑ ArcE↑ HellaS↑ PiQA↑ WinoG↑ Acc↑
- 2.62GiB 1.5B 16 8.58 12.54 41.21 75.51 50.12 75.79 63.85 61.30
GPTQ 2.45GiB 7B 3 10.73 13.90 39.33 71.21 50.16 73.29 60.77 58.95
NeUQI 2.45GiB 7B 3 8.64 11.72 44.80 76.77 55.15 77.97 70.96 65.13
- 5.55GiB 3B 16 7.44 11.15 45.22 77.31 55.03 78.62 68.75 64.99
GPTQ 4.96GiB 14B 3 8.48 11.61 38.82 67.21 54.22 75.52 65.51 60.26
NeUQI 4.96GiB 14B 3 6.82 9.89 50.17 80.43 59.39 79.54 76.09 69.12
- 13.05GiB 7B 16 6.39 10.02 48.29 80.56 60.00 78.67 72.69 68.04
GPTQ 11.72GiB 32B 3 7.21 10.36 44.03 74.96 58.75 77.97 66.22 64.38
NeUQI 11.72GiB 32B 3 5.85 9.31 51.19 80.98 62.30 80.14 75.14 69.95
- 26.42GiB 14B 16 4.93 8.75 55.80 82.37 63.38 81.07 74.66 71.46
GPTQ 26.36GiB 72B 3 6.89 9.96 46.93 78.03 61.51 78.62 71.43 67.30
NeUQI 26.36GiB 72B 3 4.99 8.37 56.40 83.75 64.92 81.83 78.45 73.07
Quantization Overhead
Table 10 presents the quantization overhead of GPTQ and NeUQI across different model sizes from the LLaMA 2 family on a single NVIDIA A40. Taking the 7B model as an example, NeUQI requires 5.85 hours with group-wise quantization and 7.10 hours with channel-wise quantization, both without coarse-to-fine search. With the strategy enabled, the overhead dropped to 0.45 and 0.50 hours, respectively, achieving over 10× speedup.
6 Conclusion
To the best of our knowledge, our proposed NeUQI is the first work to remove the zero-point integer constraint in parameter initialization for $\geq 2$-bit uniform quantization. It consistently outperforms existing methods and, under the comparable memory footprint, can even surpass non-quantized models from the same model family, improving the practicality of low-bit-width quantization. A lightweight knowledge distillation experiment further supports the effectiveness of NeUQI and our belief that combining it with other orthogonal techniques could yield additional gains. In future work, we plan to further optimize the original quantization loss during initialization and explore more efficient low-bit-width quantization methods, especially for ultra-large models.
Appendix A Limitations of Min-Max Initialization
While practitioners widely adopt min-max initialization as a matter of convention, it is not optimal concerning quantization error—even under ideal conditions. To illustrate this, consider the following lemma:
Lemma 1.
If weights are independently and uniformly drawn from an interval $[a,b]$, then the optimal parameters to for $k$-bit uniform quantization to minimize excepted quantization error $\mathbb{E}_{x\sim\mathcal{U}(a,b)}[(Q_{s,z}(x)-x)^{2}]$ are given by
$s=\frac{b-a}{2^{k}},\quad z=\frac{a}{s}+\frac{1}{2}.$ (10)
The proof of Lemma 1 is in Appendix B. This result leads to a revised understanding of the low-error interval: it was previously defined as $[zs,(2^{k}-1+z)s]$, but is now more accurately characterized as $[(z-\frac{1}{2})s,(z+2^{k}-\frac{1}{2})s]$. Motivated by this observation, we propose an improved initialization strategy, termed Min-Max+, in which the quantization parameters are given by:
$s=\frac{\text{max}(x)-\text{min}(x)}{2^{k}},\quad z=\left\lfloor\frac{\text{min}(x)}{s}+\frac{1}{2}\right\rceil.$ (11)
Interestingly, Min-Max+ closely resembles the shrink-scale technique used in MagR, where a manually defined factor $\beta$ is used to shrink the scale, and the scale is computed as $s=\beta\cdot(\text{max}(x)-\text{min}(x))/(2^{k}-1)$. In contrast, Min-Max+ implicitly sets $\beta=(2^{k}-1)/2^{k}$ and simultaneously adjusts the zero-point to better align the low-error interval with the parameter range.
Table 5: Comparison of Min-Max and Min-Max+. $\dagger$ indicates that use shrink-scale technique.
Method Min-Max Min-Max Min-Max+
scale ($s$) $\frac{\text{max}(x)-\text{min}(x)}{2^{k}-1}$ $\frac{\text{max}(x)-\text{min}(x)}{2^{k}}$ $\frac{\text{max}(x)-\text{min}(x)}{2^{k}}$
zero-point ($z$) $\left\lfloor\frac{\text{min}(x)}{s}\right\rceil$ $\left\lfloor\frac{\text{min}(x)}{s}\right\rceil$ $\left\lfloor\frac{\text{min}(x)}{s}+\frac{1}{2}\right\rceil$
Bits Group Wiki2 C4 Wiki2 C4 Wiki2 C4
3 128 6.34 7.86 6.40 7.89 6.32 7.72
3 - 8.45 9.87 7.47 8.97 7.28 8.84
2 128 26.31 23.52 50.34 58.33 13.64 13.64
2 - 6953 2592 424 167 497 135
Empirically, this simple adjustment yields measurable improvements. Under real-world settings—such as GPTQ quantization of the LLaMA 2 7B model—Min-Max+ consistently outperforms both the traditional Min-Max initialization and the variant with only shrink scaling (i.e., without zero-point shifting), achieving lower perplexity. Detailed comparisons are in Table 5.
Appendix B Proofs of Optimal Quantization Parameter and Lemma  1
We begin by relaxing the uniform quantization constraint that requires equal spacing between adjacent quantized values in the set $\mathcal{Q}$, allowing non-uniform spacing. However, the minimum error in the mean squared error objective is achieved when the spacing is uniform.
The expression for the expected quantization error takes the form
[…]
where $N=2^{k}$, $a=b_{0}<\cdots<b_{N}=b$, $q_{i}$ is the $i$-th quantized value and $[b_{i},b_{i+1})$ assumed to be quantized to $q_{i}$. The probability density function is given by $f_{X}(x)=1/(b-a)$.
For fixed interval, minimizing
[…]
requires solving
[…]
which gives
[…]
The corresponding quantization error is
[…]
Denote $\Delta_{i}=b_{i+1}-b_{i}>0$ with $\sum_{i=0}^{N-1}\Delta_{i}=b-a$. Then
$\mathcal{E}=\frac{1}{12(b-a)}\sum_{i=0}^{N-1}\Delta_{i}^{3}.$ (17)
Since $g(\Delta)=\Delta^{3}$ is strictly convex for $\Delta>0$, Jensen’s inequality implies $\sum\Delta_{i}^{3}$ is minimized when all $\Delta_{i}$ are equal, i.e.
$\Delta_{i}=\frac{b-a}{N}=\frac{b-a}{2^{k}}.$ (18)
Finally, the optimal uniform quantization parameters can be obtained by:
$s=\Delta_{i}=\frac{b-a}{2^{k}},\quad z=\frac{q^{*}_{0}}{s}=\frac{b_{0}+b_{1}}{2s}=\frac{a+(a+s)}{2s}=\frac{a}{s}+\frac{1}{2}.$ (19)
Appendix C Zero-Point Search Algorithms
Algorithm 2 Optimal Zero-Point Search over Smoothed Objective
1:Samples $\{x_{i}\}_{i=1}^{n}$, bit-width $k$
2:Optimal zero-point $z_{\mathrm{rough}}$ minimizing smoothed quantization loss
3:Initialize list of transition points: $\mathcal{T}\leftarrow[]$
4:for each sample $x_{i}$ do
5:     Compute enter transition: $t_{\mathrm{enter}}\leftarrow x_{i}-(2^{k}-0.5)$
6:     Compute enter transition loss increment: $\Delta\mathcal{L}_{\mathrm{enter}}\leftarrow h_{i}(0.25-((2^{k}-1)-(x_{i}-z))^{2})$
7:     Compute exit transition: $t_{\mathrm{exit}}\leftarrow x_{i}+0.5$
8:     Compute exit transition loss increment: $\Delta\mathcal{L}_{\mathrm{exit}}\leftarrow h_{i}((0-(x_{i}-z))^{2}-0.25)$
9:     Add $(t_{\mathrm{enter}},\Delta\mathcal{L}_{\mathrm{enter}})$ and $(t_{\mathrm{exit}},\Delta\mathcal{L}_{\mathrm{exit}})$ to $\mathcal{T}$
10:end for
11:Sort $\mathcal{T}$ by transition point $t$
12:Initialize coefficients $Az^{2}+Bz+C\leftarrow\sum_{i}h_{i}((2^{k}-1)-(x_{i}-z))^{2}$
13:Let $(t_{\text{first}},\delta_{\text{first}})\leftarrow$ the first element of $\mathcal{T}$
14:Compute minimum: $z^{\prime}\leftarrow-B/(2A)$
15:Clamp: $z^{\prime}\leftarrow\text{min}(z^{\prime},t_{\text{first}})$
16:Evaluate loss at $z^{\prime}$, record as current best: $z_{\mathrm{rough}}\leftarrow z^{\prime}$
17:for each $(t,\delta)$ in $\mathcal{T}$ do
18:     Update $Az^{2}+Bz+C$ using $\delta$
19:     Let next transition point be $t^{\prime}$ (or +$\infty$+ ∞ if none)
20:     Compute minimum: $z^{\prime}\leftarrow-B/(2A)$
21:     Clamp: $z^{\prime}\leftarrow\text{min}(\text{max}(z^{\prime},t),t^{\prime})$
22:     Evaluate loss at $z^{\prime}$, update $z_{\mathrm{rough}}$ if smaller
23:end for
24:return $z_{\mathrm{rough}}$ with minimal loss
Algorithm 3 Optimal Zero-Point Search over Piecewise Quadratic Function in Limited Interval
1:Samples $\{x_{i}\}_{i=1}^{n}$, bit-width $k$, interval $[z_{l},z_{r}]$
2:Optimal zero-point $z^{*}$ minimizing simplified quantization loss in interval $[z_{l},z_{r}]$
3:Initialize list of transition points: $\mathcal{T}\leftarrow[]$
4:for each sample $x_{i}$ do
5:     for $j=0$ to $22^{k}-$ 2 do
6:         Compute transition point: $t\leftarrow x_{i}-(j+0.5)$
7:         Compute loss increment: $\Delta\mathcal{L}\leftarrow h_{i}((j-(x_{i}-z))^{2}-(j+1-(x_{i}-z))^{2})$
8:         if $t\in[z_{l},z_{r})$ then
9:              Add $(t,\Delta\mathcal{L})$ to $\mathcal{T}$
10:         end if
11:     end for
12:end for
13:Sort $\mathcal{T}$ by transition point $t$
14:Initialize coefficients $Az^{2}+Bz+C\leftarrow\sum_{i}h_{i}(\text{clamp}(\lfloor x_{i}-z_{l}\rceil,0,2^{k}-1)-(x_{i}-z))^{2}$
15:Let $(t_{\text{first}},\delta_{\text{first}})\leftarrow$ the first element of $\mathcal{T}$
16:Compute minimum: $z^{\prime}\leftarrow-B/(2A)$
17:Clamp: $z^{\prime}\leftarrow\text{min}(\text{max}(z^{\prime},z_{l}),t_{\text{first}})$
18:Evaluate loss at $z^{\prime}$, record as current best: $z^{*}\leftarrow z^{\prime}$
19:for each $(t,\delta)$ in $\mathcal{T}$ do
20:     Update $Az^{2}+Bz+C$ using $\delta$
21:     Let next transition point be $t^{\prime}$ (or $z_{r}$ if none)
22:     Compute minimum: $z^{\prime}\leftarrow-B/(2A)$
23:     Clamp: $z^{\prime}\leftarrow\text{min}(\text{max}(z^{\prime},t),t^{\prime})$
24:     Evaluate loss at $z^{\prime}$, update $z^{*}$ if smaller
25:end for
26:return $z^{*}$ with minimal loss
Appendix D Implementation Details
D.1 Setting
D.1.1 Calibration Stage
During the calibration stage, we follow the implementation of the official GPTQ repository, [^2] https://github.com/IST-DASLab/gptq and draw the same 128 samples from the C4 dataset for quantization calibration. For implementation convenience, we uniformly adopt bfloat16 for all experiments. The token length of each sample is determined based on the characteristics and design intent of the target family. Especially for the LLaMA 2 family, each sample consists of 2048 tokens, consistent with the configurations adopted in GPTQ and LeanQuant. For the Qwen 2.5 models, each sample comprises 4096 tokens, as indicated in the Qwen 2.5 Technical Report [32], which specifies a 4096-token context length in its base configuration before context expansion. The same token length setting is applied to the LLaMA 3 family to ensure consistency.
D.1.2 Distillation Stage
In contrast to the calibration stage, the distillation utilizes 256 samples from the C4 dataset. The token lengths per sample remain consistent with those used during calibration. Furthermore, the computation precision remains bfloat16, identical to that in the calibration phase.
For the broader fine-tuning setup, we employ the AdamW optimizer with zero weight decay and momentum parameters set as $\beta=(0.9,0.95)$. The learning rate is chosen via grid search over five candidate values: 1e-5, 3e-5, 1e-4, 3e-4, and 1e-3. We apply a cosine learning rate scheduler with a 10% warm-up ratio. In terms of hardware and infrastructure, we fine-tune LLaMA 2 7B in mixed precision using DeepSpeed [^3] https://github.com/microsoft/DeepSpeed, with batch size of 1 on a single NVIDIA A40 GPU.
D.1.3 Evaluation Stage
Regarding perplexity evaluation, we follow the procedure implemented in the official GPTQ repository, specifically the datautils.py [^4] https://github.com/IST-DASLab/gptq/blob/main/datautils.py and the llama.py [^5] https://github.com/IST-DASLab/gptq/blob/main/llama.py script. Perplexity scores are computed on the validation sets of the C4 and WikiText2 datasets, using the dataset identifiers c4 and wikitext2 as specified in the original implementation.
For zero-shot accuracy evaluation, we adopt the lm-evaluation-harness [9] [^6] https://github.com/EleutherAI/lm-evaluation-harness, a standardized and extensible framework for evaluating language models across diverse tasks. Inference is conducted using vLLM [16] as the backend to ensure high throughput and memory efficiency. We evaluate five benchmark tasks: ARC-Easy, ARC-Challenge, WinoGrande, HellaSwag, and PIQA, whose corresponding identifiers in lm-evaluation-harness are arc_easy, arc_challenge, winogrande, hellaswag, and piqa, respectively.
D.2 Quantization
We adopt the quantization strategy implemented in the official GPTQ repository. Specifically, we perform sequential quantization by quantizing layers in the order of the model forward pass. Each layer receives inputs from the already quantized preceding layers rather than from the original model, ensuring that the quantization is based on realistic activation distributions that more closely resemble those encountered during inference. Furthermore, we conduct the entire quantization process under full-precision settings to ensure numerical stability. For NeUQI, the grid search over the scale hyperparameter is conducted with $T=2048$ and $T_{c}=64$.
D.3 GPTQ
We implement the GPTQ [8] method using the LDLQ [5] approach. The mathematical proof of their equivalence is in Appendix D.3.3. Moreover, LDLQ is computationally more efficient, as it requires fewer matrix operations and avoids the numerically unstable process of matrix inversion.
D.3.1 GPTQ Implementation
GPTQ builds upon the principle of Optimal Brain Surgeon (OBS) [12], designed for pruning. OBS applies a Taylor expansion of the loss function around the original weight $w$. Under the assumption that the model has reached convergence, the first-order term vanishes as a result of the stationarity condition, and higher-order terms of order three and above are considered negligible. Therefore, the dominant contribution is given by:
$\mathcal{L}(\hat{w})\approx\frac{1}{2}(\hat{w}-w)^{\top}\frac{\partial^{2}\mathcal{L}}{\partial\hat{w}^{2}}(\hat{w}-w)+\mathcal{L}(w)$ (20)
The quantization loss for a single-row in linear layer is the squared $\ell_{2}$-norm between the quantized output and the original output:
$\mathcal{L}(\hat{w})=\frac{1}{2}\|X\hat{w}-Xw\|_{2}^{2},$ (21)
where $\hat{w}$ denotes the quantized weight vector and $w$ represents the original weight vector. This formulation follows a standard least-squares loss structure. The Hessian matrix of this loss with respect to $\hat{w}$ is computed as:
$\nabla^{2}_{\hat{w}}\mathcal{L}(\hat{w})=\frac{\partial^{2}}{\partial\hat{w}^{2}}\frac{1}{2}\|X\hat{w}-Xw\|_{2}^{2}=X^{\top}X.$ (22)
Thus, the matrix $H$, representing the Hessian matrix, is given by:
$H=\nabla^{2}_{\hat{w}}\mathcal{L}(\hat{w})=X^{\top}X.$ (23)
To ensure that the Hessian matrix $H$ is positive definite, it is common practice to add a small scalar $\lambda>0$ to its diagonal entries. This modification is equivalent to augmenting the quantization loss with a regularization term $\frac{\lambda}{2}\|\hat{w}-w\|^{2}$, which penalizes large deviations between the quantized weights $\hat{w}$ and the original weights $w$. In practice, $\lambda$ is set to $0.01$ times the mean of the diagonal entries of $H$, following GPTQ.
Given this formulation, we consider the problem of quantizing a single column vector $w$ from the weight matrix. Specifically, let us focus on quantizing the element at a particular position $p$ within the vector $w$. In other words, $w_{p}$ denotes the scalar entry at the $p$-th position of $w$. We define a compensation vector $\delta^{p}$, which compensates for the quantization-induced error in $w$ resulting from quantizing $w_{p}$. In this formulation, the update at coordinate $p$ reflects the quantization of $w_{p}$, while changes at other coordinates are to compensate for the loss caused by this quantization. The quantized vector is then given by $\hat{w}=w+\delta^{p}$, with the constraint that the $p$-th coordinate of $\hat{w}$ matches the quantized value. Accordingly, we solve the following optimization problem:
$\text{min}_{\delta^{p}}$ $\quad(\delta^{p})^{\top}H\delta^{p}$ (24)
s.t. $\quad e_{p}^{\top}\delta^{p}=\hat{w}_{p}-w_{p}$
Using the method of Lagrange multipliers, we construct the Lagrangian:
$\mathcal{L}(\delta^{p},\lambda)=(\delta^{p})^{\top}H\delta^{p}+\lambda\left(e_{p}^{\top}\delta^{p}-(\hat{w}_{p}-w_{p})\right)$ (25)
Taking the gradient with respect to $\delta^{p}$ and setting it to zero:
$\nabla_{\delta^{p}}\mathcal{L}=2H\delta^{p}+\lambda e_{p}=0\quad\Rightarrow\quad\delta^{p}=-\frac{\lambda}{2}H^{-1}e_{p}$ (26)
Substituting back into the constraint:
$e_{p}^{\top}\delta^{p}=-\frac{\lambda}{2}e_{p}^{\top}H^{-1}e_{p}=\hat{w}_{p}-w_{p}\quad\Rightarrow\quad\lambda=-\frac{2(\hat{w}_{p}-w_{p})}{(H^{-1})_{pp}}$ (27)
Thus, we obtain:
$\delta^{p}=\frac{(\hat{w}_{p}-w_{p})}{(H^{-1})_{pp}}H^{-1}e_{p},\quad\mathcal{L}(\delta^{p})=\frac{(\hat{w}_{p}-w_{p})^{2}}{(H^{-1})_{pp}}$ (28)
Based on this result, we can adopt a greedy strategy: at each iteration, the position $p$ that results in the minimal increase in quantization loss $\mathcal{L}(\delta^{p})$ is selected, the corresponding update is applied. However, to better utilize GPU parallelism and improve computational efficiency, we follow a heuristic implemented in the GPTQ repository, specifically in the gptq.py script  [^7] https://github.com/IST-DASLab/gptq/blob/main/gptq.py. In this approach, the indices are sorted and traversed in decreasing order of the Hessian diagonals $H_{pp}$.
Assuming we fix a quantization order $i=0,1,\ldots,n-1$, we define:
$\delta^{i}_{i:}=\frac{(H_{i:,i:})^{-1}e_{i}}{((H_{i:,i:})^{-1})_{ii}},\quad\delta^{i}_{:i}=\textbf{0}$ (29)
where, $\delta^{i}$ denotes the update basis vector for position $i$, satisfying $(\delta^{i})_{i}=1$. The actual update applied to the weight vector is $(\hat{w}_{i}-w_{i})\delta^{i}$.
For any $j\geq i$, we have:
$H_{j,:}\delta^{i}=e_{j}^{\top}H_{i:,i:}\left(\frac{(H_{i:,i:})^{-1}e_{i}}{((H_{i:,i:})^{-1})_{ii}}\right)=1_{i=j}\left(((H_{i:,i:})^{-1})_{ii}\right)^{-1}$ (30)
Therefore, the update matrix satisfies:
$H\Delta=U,$ (31)
where $\Delta=[\delta^{0},\delta^{1},\ldots,\delta^{n-1}]$ is a unit lower-triangular matrix, and the resulting matrix $U$ is upper-triangular because for $j>i$, $U_{j,i}=0$.
This relation implies that $H$ can be factorized as
$H=U\Delta^{-1},$ (32)
which corresponds to a UL decomposition of $H$, where $\Delta^{-1}$ is unit lower-triangular. On the other hand, the Hessian matrix $H$ is positive-definite and thus admits a unique LDL decomposition of the form
$H=(L^{\top}D)L,$ (33)
where $L$ is also a unit lower-triangular matrix and $D$ is diagonal matrix.
Since both (32) and (33) represent decompositions of the same matrix $H$ with a unit lower-triangular factor on the right, and the unit lower-triangular factor in such factorizations is unique for positive-definite matrices, we conclude that $\Delta=L^{-1}$.
Finally, the algorithm is as follows:
Algorithm 4 GPTQ Quantization
1:Hessian matrix $H$, original weight vector $w$, quantization operator $Q(\cdot)$
2:Compute LDL decomposition: $H=L^{\top}DL$
3:Compute inverse of unit lower-triangular matrix: $\Delta=L^{-1}$
4:Compute compensation matrix: $L\leftarrow\Delta-I$
5:Initialize quantized weight $\hat{w}$ with zeros
6:for each index $i=0$ to $n-1$ do
7:     $\hat{w}_{i}\leftarrow Q(w_{i})$
8:     $w\leftarrow w+L_{:,i}\cdot(\hat{w}_{i}-w_{i})$
9:end for
10:return $\hat{w}$
D.3.2 LDLQ Implementation
LDLQ is a newly proposed method introduced in QuIP [5], which has been empirically validated to be equivalent to GPTQ. It provides an alternative understanding of the GPTQ algorithm through a structurally distinct formulation. Specifically, when quantizing a single row of the weight matrix, LDLQ proceeds coordinate-wise in a fixed order: each coordinate is quantized one at a time, and the unquantized coordinates are adjusted to compensate for the quantization error. This process is captured through a strictly lower-triangular compensation matrix $L$, which defines the final quantized weights via
$\hat{w}=Q(\tilde{w}),\quad\tilde{w}=w-L(\hat{w}-w)$ (34)
The quantization-induced error is
$\eta=\hat{w}-\tilde{w}=(I+L)(\hat{w}-w)$ (35)
Solving for $\hat{w}-w$ yields
$\hat{w}-w=(I+L)^{-1}\eta$ (36)
Substituting into the original quadratic loss function gives
$\mathcal{L}(\hat{w})=\eta^{\top}(I+L)^{-\top}H(I+L)^{-1}\eta$ (37)
Given the LDL decomposition $H=L^{\prime\top}DL^{\prime}$, LDLQ claims that the minimal loss is achieved when $I+L=L^{\prime}$, resulting in
$\mathcal{L}(\hat{w})=\eta^{\top}D\eta$ (38)
This diagonalization simplifies the loss into a weighted squared norm of the quantization error, enabling an efficient sequential update procedure as described below:
Algorithm 5 LDLQ Quantization
1:Hessian matrix $H$, original weight vector $w$, quantization operator $Q(\cdot)$
2:Compute LDL decomposition: $H=L^{\prime\top}DL^{\prime}$
3:Initialize quantized weight vector $\hat{w}$ with zeros
4:Compute compensation matrix $L\leftarrow L^{\prime}-I$
5:for each index $i=0$ to $n-1$ do
6:     $\tilde{w}_{i}\leftarrow w_{i}-L_{i,:}\cdot(\hat{w_{i}}-w_{i})$
7:     $\hat{w}_{i}\leftarrow Q(\tilde{w}_{i})$
8:end for
9:return $\hat{w}$
D.3.3 Equivalence between GPTQ and LDLQ
The following interpretation reinforces the equivalence between GPTQ and LDLQ. In particular, we can rewrite the GPTQ formulation as follows:
$\hat{w}=Q(\tilde{w}),\quad\tilde{w}=w+(\Delta-I)(\hat{w}-\tilde{w})$ (39)
We can derive a more explicit form of $\tilde{w}$ by manipulating the second equation in (39) as follows:
$\tilde{w}$ $=w+(\Delta-I)(\hat{w}-\tilde{w})$ (40)
$\Delta\tilde{w}$ $=w+(\Delta-I)\hat{w}$ (41)
$\tilde{w}$ $=\Delta^{-1}w+(I-\Delta^{-1})\hat{w}$ (42)
$\tilde{w}$ $=w-(\Delta^{-1}-I)(\hat{w}-w)$ (43)
On the other hand, the LDLQ formulation is given by:
$\tilde{w}=w-(L^{\prime}-I)(\hat{w}-w)$ (44)
By comparing equations (43) and (44), we observe that the two formulations become equivalent if $\Delta^{-1}=L^{\prime}$. As established in the previous two sections, both $\Delta^{-1}$ and $L^{\prime}$ correspond to the $L$ factor derived from the LDL decomposition. Therefore, LDLQ is mathematically equivalent to GPTQ.
D.4 MagR
For MagR [34], we strictly follow the original settings, including using the hyperparameter $a$ of 1e-3 for channel-wise quantization and 1e-4 for group-wise quantization, as well as adopting the shrink-scale technique with the same $\beta$ value as in the original paper.
D.5 Difference between LeanQuant and NeUQI with Integer Constraint
[…]
$\mathcal{L}(s,z)=\sum_{i}h_{i}(Q_{s,z}(w_{i})-w_{i})^{2},$ (45)
but differ fundamentally in how the weight importance $h_{i}$ is defined and how to find the optimal scale and zero-point.
[…]
$\frac{(Q_{s,z}(w_{i})-w_{i})^{2}}{(H^{-1})_{ii}},$ (46)
leading to $h_{i}=((H^{-1})_{ii})^{-1}$, and further generalized to $h_{i}=((H^{-1})_{ii})^{-p}$ with a tunable strength parameter $p$. In contrast, $\text{NeUQI}_{\text{int}}$ adopts a diagonal loss approximation by directly setting $h_{i}=H_{ii}$, avoiding the need to invert the Hessian matrix.
[…]
D.6 PV-tuning
[…]
Appendix E Broader Impact
Large language models are inherently dual-use technologies, enabling beneficial applications and potential misuse. While our method does not enhance model capabilities, it facilitates broader deployment by reducing quantization-induced errors and improving consistency with the original model. This contributes to more reliable and trustworthy outputs, potentially lowering the risk of unintended behaviors such as hallucinations or misleading responses.
Appendix F Limitations
Although we do not compare NeUQI with other quantization methods based on weight and activation transformations or fine-tuning, we believe our main experiments are sufficient to validate the effectiveness of our approach. These alternative methods typically require additional training or model modifications and adopt fundamentally different design principles. As such, they are orthogonal to our NeUQI and could be integrated with NeUQI to achieve further improvements, which we leave as an avenue for future work. Our scheme supports arbitrary zero-points, which may not align with inference kernels such as ExLlamaV2888https://github.com/turboderp-org/exllamav2. Nonetheless, NeUQI is fully compatible with flexible backends like FLUTE [11], LUT-GEMM [24], and BitBLAS [30], which support mixed precision or lookup-based matrix multiplication.
Appendix G Supplementary Results
This appendix presents supplementary quantization results. Detailed results on 2-bit quantization with group size 128 for LLaMA and Qwen families are shown in Table 6. Detailed results on 3-bit and 4-bit quantization settings for LLaMA 2, Qwen 2.5 and LLaMA 3 families are shown in Table 7, Table 8 and Table 9, respectively. Quantization overhead for LLaMA 2 family are shown in Table 10.
Table 6: Results of 2-bit group-wise quantization with group size of 128 on the LLaMA 2, LLaMA 3, and Qwen 2.5 families.
Size Bits Method Wiki2↓ C4↓ ArcC↑ ArcE↑ HellaS↑ PiQA↑ WinoG↑ Acc↑
LLaMA 2 Family
7B 2 GPTQ 26.31 23.52 22.70 36.91 34.39 60.50 54.62 41.82
MagR 15.41 15.59 24.66 47.90 39.19 65.89 58.80 47.29
NeUQI 12.35 13.48 25.26 59.72 40.71 69.64 61.33 51.33
13B 2 GPTQ 12.50 13.16 26.54 53.58 42.00 68.34 55.01 49.09
MagR 17.15 11.35 30.46 61.74 45.72 71.60 60.62 54.03
NeUQI 8.82 10.60 35.32 67.72 45.98 73.12 68.59 58.15
70B 2 GPTQ 7.04 8.62 38.48 70.20 53.00 75.14 69.93 61.35
MagR 10.87 9.21 41.89 68.64 55.13 68.50 73.64 61.56
NeUQI 5.60 7.51 47.53 77.86 56.90 78.24 75.53 67.21
LLaMA 3 Family
8B 2 GPTQ 210 121 19.54 28.91 26.94 54.73 51.62 36.35
MagR 61.71 74.78 18.43 33.80 28.72 56.69 51.70 37.87
NeUQI 28.90 26.83 23.29 50.04 36.04 63.38 59.91 46.53
70B 2 GPTQ 26.25 28.59 22.27 38.80 34.61 61.04 55.41 42.43
MagR 30.86 55.42 26.71 51.56 40.78 67.74 64.56 50.27
NeUQI 10.75 14.40 39.68 71.30 48.66 74.05 70.56 60.85
Qwen 2.5 Family
7B 2 GPTQ 21.66 22.77 23.72 43.43 39.32 62.02 53.75 44.45
MagR 13.65 16.56 30.46 61.95 43.88 68.50 60.30 53.02
NeUQI 14.43 16.65 36.26 67.76 45.56 71.55 65.04 57.23
14B 2 GPTQ 18.62 19.22 22.95 42.80 38.38 61.97 51.62 43.54
MagR 12.94 14.09 28.33 56.57 44.52 70.40 60.77 52.12
NeUQI 10.57 13.61 43.52 75.63 50.26 74.92 72.30 63.32
32B 2 GPTQ 10.95 13.36 30.38 59.30 47.56 71.38 56.59 53.04
MagR 8.82 11.65 36.69 69.57 52.23 75.57 64.40 59.69
NeUQI 8.51 11.41 46.08 76.39 55.86 78.02 75.14 66.30
72B 2 GPTQ 10.35 12.27 35.49 63.51 50.11 72.36 59.35 56.17
MagR 11.00 11.22 44.37 75.88 56.58 77.31 71.59 65.15
NeUQI 6.63 9.97 51.88 81.61 59.23 79.87 75.69 69.66
Table 7: Results for LLaMA 2 family models at bfloat16 precision, as well as for the 3-bit, 3-bit with a group size of 128, 4-bit, and 4-bit with a group size of 128 quantized models.
Size Bits Group Method Wiki2↓ C4↓ ArcC↑ ArcE↑ HellaS↑ PiQA↑ WinoG↑ Acc↑
7B 16 - - 5.12 6.63 43.34 76.26 57.17 77.97 68.98 64.74
4 128 GPTQ 5.62 7.12 43.52 75.55 56.27 77.48 69.85 64.53
NeUQI 5.60 7.09 42.92 75.17 56.19 77.64 69.46 64.27
- GPTQ 5.84 7.36 41.64 74.16 55.84 77.69 69.85 63.84
NeUQI 5.80 7.26 41.89 74.87 54.78 76.77 68.51 63.36
3 128 GPTQ 6.34 7.86 39.76 73.74 53.91 77.31 67.48 62.44
NeUQI 6.07 7.58 38.57 72.35 54.19 76.44 67.48 61.81
- GPTQ 8.45 9.87 34.73 66.46 49.08 73.34 64.40 57.60
NeUQI 6.56 8.10 37.97 71.42 50.75 75.41 68.19 60.75
13B 16 - - 4.57 6.05 48.21 79.46 60.09 79.11 72.30 67.83
4 128 GPTQ 5.00 6.56 47.70 78.28 59.61 78.62 72.53 67.35
NeUQI 4.98 6.56 47.10 79.21 59.55 79.11 72.30 67.45
- GPTQ 5.15 6.71 44.97 77.44 58.88 77.91 71.27 66.09
NeUQI 5.09 6.67 45.31 77.86 58.64 78.13 71.19 66.23
3 128 GPTQ 5.43 7.05 45.22 77.23 58.14 77.69 70.72 65.80
NeUQI 5.32 6.91 45.82 78.32 57.79 78.40 72.77 66.62
- GPTQ 6.46 8.03 38.91 73.48 55.18 76.39 68.35 62.46
NeUQI 5.70 7.25 42.75 75.84 56.06 77.64 70.72 64.60
70B 16 - - 3.12 4.97 54.52 82.66 64.76 82.15 77.43 72.30
4 128 GPTQ 3.42 5.58 54.69 82.74 64.45 81.83 77.19 72.18
NeUQI 3.41 5.58 54.95 82.83 64.29 82.05 77.98 72.42
- GPTQ 3.59 5.68 54.27 81.99 64.16 82.15 77.19 71.95
NeUQI 3.47 5.62 54.35 83.08 64.04 81.66 78.53 72.33
3 128 GPTQ 3.87 5.86 53.24 81.57 63.16 81.61 77.43 71.40
NeUQI 3.71 5.77 54.95 82.45 62.99 81.99 77.03 71.88
- GPTQ 4.83 6.57 49.15 79.38 60.60 80.36 74.35 68.77
NeUQI 3.90 5.90 53.50 83.00 62.58 81.61 76.72 71.48
Table 8: Results for LLaMA 3 family models at bfloat16 precision, as well as for the 3-bit, 3-bit with a group size of 128, 4-bit, and 4-bit with a group size of 128 quantized models.
Size Bits Group Method Wiki2↓ C4↓ ArcC↑ ArcE↑ HellaS↑ PiQA↑ WinoG↑ Acc↑
8B 16 - - 5.76 8.32 50.34 80.22 60.19 79.60 73.64 68.80
4 128 GPTQ 6.19 8.99 47.61 77.86 59.06 77.75 73.88 67.23
NeUQI 6.12 8.87 50.00 80.18 59.53 78.67 74.90 68.66
- GPTQ 6.97 9.95 44.54 77.27 57.87 77.15 73.32 66.03
NeUQI 6.67 9.41 47.10 77.69 58.13 79.54 74.59 67.41
3 128 GPTQ 8.30 11.50 40.19 71.84 54.43 76.22 70.72 62.68
NeUQI 7.45 10.49 46.50 79.34 55.28 77.69 73.48 66.46
- GPTQ 19.03 29.26 25.34 46.25 43.02 62.84 59.91 47.47
NeUQI 9.70 11.61 43.77 76.30 53.87 77.20 71.35 64.50
70B 16 - - 2.68 5.88 60.49 86.95 66.37 82.48 80.90 75.44
4 128 GPTQ 3.40 6.41 57.76 85.14 66.00 82.05 80.03 74.20
NeUQI 3.17 6.25 58.02 84.76 65.82 82.43 80.90 74.39
- GPTQ 1486 1404 19.97 25.21 30.47 52.94 50.83 35.88
NeUQI 4.90 10.00 51.96 80.47 62.29 79.49 63.61 67.57
3 128 GPTQ 5.30 8.33 51.37 80.93 62.73 80.90 75.53 70.29
NeUQI 4.63 7.58 54.52 83.42 62.50 80.47 80.03 72.19
- GPTQ 2645 $1$ 20.82 25.04 26.16 52.18 51.22 35.08
NeUQI 9.04 13.36 35.84 69.11 52.97 71.82 55.80 57.11
Table 9: Results for Qwen 2.5 family models at bfloat16 precision, as well as for the 3-bit, 3-bit with a group size of 128, 4-bit, and 4-bit with a group size of 128 quantized models.
Size Bits Group Method Wiki2↓ C4↓ ArcC↑ ArcE↑ HellaS↑ PiQA↑ WinoG↑ Acc↑
7B 16 - - 6.39 10.02 48.29 80.56 60.00 78.67 72.69 68.04
4 128 GPTQ 6.61 10.19 48.21 80.35 59.24 79.05 71.35 67.64
NeUQI 6.55 10.17 48.63 80.98 59.36 78.67 73.95 68.32
- GPTQ 7.06 10.61 46.59 79.25 58.08 78.56 68.59 66.21
NeUQI 6.82 10.42 48.55 79.38 58.63 78.78 71.90 67.45
3 128 GPTQ 7.42 10.94 43.77 75.63 56.71 77.37 65.75 63.85
NeUQI 7.14 10.80 46.42 80.05 56.46 78.35 70.40 66.34
- GPTQ 10.73 13.90 39.33 71.21 50.16 73.29 60.77 58.95
NeUQI 8.64 11.72 44.80 76.77 55.15 77.97 70.96 65.13
14B 16 - - 4.93 8.75 55.80 82.37 63.38 81.07 74.66 71.46
4 128 GPTQ 5.29 8.93 54.35 81.94 62.94 80.96 74.98 71.03
NeUQI 5.20 8.89 56.40 83.25 62.78 81.01 75.61 71.81
- GPTQ 5.80 9.24 52.73 81.44 62.27 80.20 74.98 70.32
NeUQI 5.46 9.05 52.56 81.10 62.62 80.36 77.19 70.77
3 128 GPTQ 6.29 9.63 48.55 79.38 60.06 79.16 72.14 67.86
NeUQI 5.91 9.42 52.90 81.02 60.71 80.90 76.72 70.45
- GPTQ 8.48 11.61 38.82 67.21 54.22 75.52 65.51 60.26
NeUQI 6.82 9.89 50.17 80.43 59.39 79.54 76.09 69.12
32B 16 - - 4.67 8.59 53.16 80.51 65.00 81.99 75.69 71.27
4 128 GPTQ 4.87 8.69 54.10 81.44 64.59 80.69 75.93 71.35
NeUQI 4.82 8.68 53.24 80.26 64.71 81.18 76.64 71.20
- GPTQ 5.23 8.89 51.19 79.29 63.96 80.41 75.45 70.06
NeUQI 4.98 8.77 53.07 81.48 64.46 81.77 77.35 71.63
3 128 GPTQ 5.63 9.11 51.11 79.50 62.83 80.36 75.14 69.79
NeUQI 5.30 8.99 51.71 82.07 62.92 80.90 76.40 70.80
- GPTQ 7.21 10.36 44.03 74.96 58.75 77.97 66.22 64.38
NeUQI 5.85 9.31 51.19 80.98 62.30 80.14 75.14 69.95
72B 16 - - 3.64 7.75 58.11 84.76 67.59 82.10 77.35 73.98
4 128 GPTQ 3.82 7.85 57.51 84.60 67.14 82.26 78.06 73.91
NeUQI 3.78 7.83 57.51 84.47 67.16 81.99 78.93 74.01
- GPTQ 4.17 8.35 55.38 84.68 66.68 81.66 77.90 73.26
NeUQI 3.95 7.91 58.02 84.34 66.81 82.15 77.90 73.85
3 128 GPTQ 4.55 8.53 55.97 83.92 65.66 80.79 77.74 72.82
NeUQI 4.26 8.11 54.52 83.75 65.85 82.21 78.06 72.88
- GPTQ 6.89 9.96 46.93 78.03 61.51 78.62 71.43 67.30
NeUQI 4.99 8.37 56.40 83.75 64.92 81.83 78.45 73.07
Table 10: Quantization overhead (in hours) for LLaMA 2 family under different configurations in a single NVIDIA A40.
Size 7B 7B 13B 13B 70B 70B
Group 128 - 128 - 128 -
GPTQ 0.21 0.15 0.30 0.26 1.42 1.28
NeUQI 0.45 0.50 0.81 0.88 4.04 4.88
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: For the effectiveness of the NeUQI method, see Section 5.3 and Section 5.4. For the limitations introduced by Min-Max initialization, see the paragraph Integer Constraint in Section 5.4, as well as the ideal-case analysis in Appendix A.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: In Section F, we discuss the limitations.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: The proof for the optimal parameters of the quantized uniform distribution is provided in Appendix B, and the equivalence between GPTQ and LDLQ is discussed in Appendix D.3.3.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We quantize open-access LLMs on open datasets and release our full code. Moreover, we make every effort to document all implementation details in Appendix D, and provide instructions and hyperparameters within the code.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: As we state above, we release the full implementation for the NeUQI method with requisite instructions. We do not introduce new datasets or models, and rely solely on openly available ones.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We describe the necessary experimental settings in the implementation details of Section 5, provide a complete account of technical details in Appendix D, and include more extensive and concrete information in the supplementary code.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [N/A]
  34. Justification: For ease of comparison, we follow prior work by using a fixed random seed, and therefore do not report error bars.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We report the hardware setup and quantization overhead in the paragraph Quantization Overhead of Section 5.4, and detail the distillation stage in Appendix D.1.2. We omit details that have negligible impact on quantization overhead.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Our research is focused on the base capability and accessibility of LLMs. While working on LLMs always has potential externalities, our specific work adheres to the ethics guidelines.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: While our work is more concerned with fundamental matters of quantization parameter optimization in initialization stage, we provide a brief overview of its societal impacts in Appendix E.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: Our work does not involve the release of any new models. Moreover, quantizing existing models generally leads to a reduction in model capacity, which in turn lowers the associated risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We use academically published artifacts (datasets, models, etc.), follow their respective licenses, and cite the original authors accordingly.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We release the code and provide detailed guidance in the form of README. Both are included in the supplementary archive.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: We do not use human subjects in our experiments.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: We did not conduct any research on human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: We only use LLMs for editing (e.g., grammar, spelling, word choice).
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.